# XAI-Driven Skin Disease Classification: Leveraging GANs to Augment ResNet-50 Performance

### v1

@@ Abstract
Accurate and timely diagnosis of multi-class skin lesions is hampered by subjective methods, inherent data imbalance in datasets like HAM10000, and the ”black box” nature of Deep Learning (DL) models. This study proposes a trustworthy and highly accurate Computer-Aided Diagnosis (CAD) system to overcome these limitations. The approach utilizes Deep Convolutional Generative Adversarial Networks (DCGANs) for per-class data augmentation to resolve the critical class imbalance problem. A fine-tuned ResNet-50 classifier is then trained on the augmented dataset to classify seven skin disease categories. Crucially, LIME and SHAP Explainable AI (XAI) techniques are integrated to provide transparency by confirming that predictions are based on clinically relevant features like irregular morphology. The system achieved a high overall Accuracy of 92.50% and a Macro-AUC of 98.82%, successfully outperforming various prior benchmarked architectures. This work successfully validates a verifiable framework that combines high performance with the essential clinical interpretability required for safe diagnostic deployment. Future research should prioritize enhancing discrimination for critical categories, such as Melanoma_NOS (F11-Score: 0.8602).
@@ I Introduction and Motivation
Skin diseases and cancer are major public health concern, with cancer being a leading cause of death worldwide and skin cancer ranking as the fifth most prevalent type globally [10455953]. The accurate and early diagnosis of skin lesions is vital for effective treatment and improved patient outcomes, especially given that the highly aggressive forms like Melanoma have a 99% survival rate when detected early [8807441, 10880715]. However, traditional diagnostic methods relying on subjective visual inspection, expert judgment, and often expensive and invasive biopsies can be time-consuming, inconsistent, and less accurate, particularly in regions with limited access to specialists [10020302, 11210939, 11052018, 10924047, 10455953].
@@
This study addresses these two challenges by proposing a robust and trustworthy automated skin lesion classification system specifically focusing on data augmentation using Generative Adversarial Networks (GANs) to synthesize high quality images for minority classes, overcoming the class imbalance and data diversity; usage of CNN model Resnet50 architecture, a powerful proven CNN framework, for classifying augmented dataset into multiple skin disease categories, and enhancing clinical trust with Explainable AI to provide transparency and make the system clinically reliable highlighting the image features driving the model’s predictions. By integrating all of these, the study aims to provide a reliable, efficient, and clinically interpretable CAD tool to significantly improve the accuracy and adoption of automated skin diseases diagnosis.
@@ III-A Data collection
The study used an extended version of the HAM10000 dataset from the ISIC Archive [tschandl2018ham10000], a publicly available multi-source repository of dermatoscopic images representing seven types of pigmented skin lesions: melanocytic nevi, keratosis subtypes, basal cell carcinoma, squamous cell carcinoma, dermatofibroma, and melanoma. The images were originally collected by the ViDIR Group at the Medical University of Vienna and a dermatology practice associated with the University of Queensland, Australia. While the canonical HAM10000 dataset contains 10,015 images, the version used in this study included 11,540 images with valid metadata. These images were used to create stratified training, validation, and test sets, ensuring balanced representation across lesion classes.
@@ III-C Synthetic Data Generation
To address class imbalance, a per-class Deep Convolutional Generative Adversarial Network (DCGAN) strategy was applied exclusively to the minority classes, allowing the model to learn the unique features of underrepresented lesions without interference from the majority class (Nevus). Generative Adversarial Networks (GANs), a prominent framework for generative modeling, utilize a competitive training setup between two neural networks—a Generator and a Discriminator—to produce highly realistic synthetic data [11020485, geeksforgeeksGenerativeAdversarial]. Fig. 3 illustrates the competitive training process between the Generator (G) and Discriminator (D). The DCGAN architecture specifically implements both G and D using convolutional layers (transpose convolutions for G and strided convolutions for D).
@@ III-D CNN model Resnet50
Model initialization follows the transfer learning approach: the original convolutional layers of ResNet-50 are frozen by setting their parameters to non-trainable, and only a new custom classification head is trained. The model is trained using the Adam optimizer with a learning rate of 0.00001 to gently update the new layers without altering the frozen weights. Training uses a batch size of 64 and employs early stopping with a patience of 5 epochs based on validation loss to prevent overfitting. The model was trained for 30 epochs. After training, performance is evaluated on the unseen test set.
@@
After training, the model’s performance was rigorously evaluated on the unseen test set using standard metrics, including Accuracy, Precision, Recall, F1-Score, and Macro-AUC (One-vs-Rest). This evaluation was essential for assessing the model’s effectiveness and its ability to generalize across different skin disease types. During the training process, we monitored both the training and validation accuracy to understand how the model progressed on both seen and unseen data.
@@
Macro-AUC (OVR): This metric (Area Under the Curve, One-vs-Rest) measures the model’s overall ability to correctly distinguish one disease class from all others, averaged across every disease type. The high percentage indicates that the model is excellent at separating all the different skin conditions across all possible decision thresholds, consistently performing much better than random chance [10924047].
@@ III-E Class-Wise Performance Analysis
The performance metrics were also stratified according to the individual skin lesion categories. The multi-class performance by the GAN-based **Resnet50** model was assessed independently across the seven defined classes, with the specific Precision, Recall, and F11-Score for each class displayed in Table III that presents class-wise performance metrics.
@@
The model demonstrated exceptional performance for **Solar or actinic keratosis** (F11-Score 0.9922) and **Dermatofibroma** (F11-Score 0.9889), achieving a perfect Precision of 1.0000 for both, meaning no false positives were recorded for these two diseases. Classification of Squamous_cell_carcinoma_NOS was also very strong (F11-Score 0.9755). However, the lowest F11-Scores were observed for Nevus (0.8541) and Melanoma_NOS (0.8602), indicating that these remain the primary classification challenges. Notably, the Melanoma_NOS class exhibited a significant trade-off, showing high Recall (0.9113)—minimizing false negatives by successfully identifying most actual cases—but lower Precision (0.8145). These result highlight the need for future efforts to improve discrimination among these more challenging skin diseases conditions.
@@ III-F Comparison with Previous Studies
The findings from the present study were compared against prior works that utilized the same dataset. Using the comparison framework presented in [10455953], and incorporating the results from the three most recent studies [10455953, 10020302, Chaturvedi2020], this study’s performance metrics were added to provide a comprehensive benchmark.
@@ IV-A Visual Validation using LIME and SHAP
To validate the model’s predictive focus, LIME and **SHAP** were applied to a representative image classified as Melanoma_NOS.
@@
Figure 9: Lime-based visualization

### v2

@@ Abstract
Accurate and timely diagnosis of multi-class skin lesions is hampered by subjective methods, inherent data imbalance in datasets like HAM10000, and the ”black box” nature of Deep Learning (DL) models. This study proposes a trustworthy and highly accurate Computer-Aided Diagnosis (CAD) system to overcome these limitations. The approach utilizes Deep Convolutional Generative Adversarial Networks (DCGANs) for per-class data augmentation to resolve the critical class imbalance problem. A fine-tuned ResNet-50 classifier is then trained on the augmented dataset to classify seven skin disease categories. Crucially, LIME and SHAP Explainable AI (XAI) techniques are integrated to provide transparency by confirming that predictions are based on clinically relevant features like irregular morphology. The system achieved a high overall Accuracy of 92.50% and an Area Under the Curve (AUC) of 98.82%, successfully outperforming various prior benchmarked architectures. This work successfully validates a verifiable framework that combines high performance with the essential clinical interpretability required for safe diagnostic deployment. Future research should prioritize enhancing discrimination for critical categories, such as Melanoma_NOS (F11-Score: 0.8602).
@@ I Introduction and Motivation
Skin diseases and cancer are major public health concern, with cancer being a leading cause of death worldwide and skin cancer ranking as the fifth most prevalent type globally [1]. The accurate and early diagnosis of skin lesions is vital for effective treatment and improved patient outcomes, especially given that the highly aggressive forms like Melanoma have a $99\%$ survival rate when detected early [2, 3]. However, traditional diagnostic methods relying on subjective visual inspection, expert judgment, and often expensive and invasive biopsies can be time-consuming, inconsistent, and less accurate, particularly in regions with limited access to specialists [4, 5, 6, 7, 1].
@@
This study addresses these two challenges by proposing a robust and trustworthy automated skin lesion classification system specifically focusing on data augmentation using Generative Adversarial Networks (GANs) to synthesize high quality images for minority classes, overcoming the class imbalance and data diversity; usage of Convolutional Neural Network (CNN) model Resnet50 architecture, a powerful proven CNN framework, for classifying augmented dataset into multiple skin disease categories, and enhancing clinical trust with Explainable AI to provide transparency and make the system clinically reliable highlighting the image features driving the model’s predictions. By integrating all of these, the study aims to provide a reliable, efficient, and clinically interpretable CAD tool to significantly improve the accuracy and adoption of automated skin diseases diagnosis.
@@ III-A Data collection
The study used an extended version of the HAM10000 dataset from the ISIC Archive [22], a publicly available multi-source repository of dermatoscopic images representing seven types of pigmented skin lesions: melanocytic nevi, keratosis subtypes, basal cell carcinoma, squamous cell carcinoma, dermatofibroma, and melanoma. The images were originally collected by the ViDIR Group at the Medical University of Vienna and a dermatology practice associated with the University of Queensland, Australia. While the canonical HAM10000 dataset contains 10,015 images, the version used in this study included 11,540 images with valid metadata. These images were used to create stratified training, validation, and test sets, ensuring balanced representation across lesion classes.
@@ III-C Synthetic Data Generation
To address class imbalance, a per-class Deep Convolutional Generative Adversarial Network (DCGAN) strategy was applied exclusively to the minority classes, allowing the model to learn the unique features of underrepresented lesions without interference from the majority class (Nevus). Generative Adversarial Networks (GANs), a prominent framework for generative modeling, utilize a competitive training setup between two neural networks—a Generator and a Discriminator—to produce highly realistic synthetic data [18, 23]. Fig. 3 illustrates the competitive training process between the Generator (G) and Discriminator (D). The DCGAN architecture specifically implements both G and D using convolutional layers (transpose convolutions for G and strided convolutions for D).
@@ III-D CNN model Resnet-50
Model initialization follows the transfer learning approach: the original convolutional layers of ResNet-50 are frozen by setting their parameters to non-trainable, and only a new custom classification head is trained. The model is trained using the Adaptive moment estimation optimizer with a learning rate of 0.00001 to gently update the new layers without altering the frozen weights. Training uses a batch size of 64 and employs early stopping with a patience of 5 epochs based on validation loss to prevent overfitting. The model was trained for 30 epochs. After training, performance is evaluated on the unseen test set.
@@
After training, the model’s performance was rigorously evaluated on the unseen test set using standard metrics, including Accuracy, Precision, Recall, F1-Score, and AUC (One-vs-Rest). This evaluation was essential for assessing the model’s effectiveness and its ability to generalize across different skin disease types. During the training process, we monitored both the training and validation accuracy to understand how the model progressed on both seen and unseen data.
@@
AUC (OVR): This metric (Area Under the Curve, One-vs-Rest) measures the model’s overall ability to correctly distinguish one disease class from all others, averaged across every disease type. The high percentage indicates that the model is excellent at separating all the different skin conditions across all possible decision thresholds, consistently performing much better than random chance [7].
@@ III-E Class-Wise Performance Analysis
The performance metrics were also stratified according to the individual skin lesion categories. The multi-class performance by the GAN-based Resnet50 model was assessed independently across the seven defined classes, with the specific Precision, Recall, and F11-Score for each class displayed in Table III that presents class-wise performance metrics.
@@
The model demonstrated exceptional performance for Solar or actinic keratosis (F11-Score 0.9922) and Dermatofibroma (F11-Score 0.9889), achieving a perfect Precision of 1.0000 for both, meaning no false positives were recorded for these two diseases. Classification of Squamous_cell_carcinoma_NOS was also very strong (F11-Score 0.9755). However, the lowest F11-Scores were observed for Nevus (0.8541) and Melanoma_NOS (0.8602), indicating that these remain the primary classification challenges. Notably, the Melanoma_NOS class exhibited a significant trade-off, showing high Recall (0.9113)—minimizing false negatives by successfully identifying most actual cases—but lower Precision (0.8145). These result highlight the need for future efforts to improve discrimination among these more challenging skin diseases conditions.
@@ III-F Comparison with Previous Studies
The findings from the present study were compared against prior works that utilized the same dataset. Using the comparison framework presented in [1], and incorporating the results from the three most recent studies [1, 4, 24], this study’s performance metrics were added to provide a comprehensive benchmark.
@@ IV-A Visual Validation using LIME and SHAP
To validate the model’s predictive focus, LIME and SHAP were applied to a representative image classified as Melanoma_NOS.
@@
Figure 9: Lime-based visualization of prediction explanations.
@@ VI Acknowledgments
VII Funding
This work was not supported by any funding agency or grant.